Calculate the confidence interval to estimate the population proportion, given the following data: a. n=400; p ̂ = 0.70; confidence level : 95% and b. n=700; p ̂ = 0.45; confidence level : 99%. 


1
Expert's answer
2021-05-12T02:04:04-0400

a.) "p = 0.70"


"n = 400"


We know confidence interval can be given as,


"CI = p \pm z \sqrt{\dfrac{p(1-p)}{n}}"


"z= 1.96" for 95% Confidence level.


"CI = 0.70 \pm 1.96 \sqrt{\dfrac{0.70\times 0.30}{400}}"


"CI = 0.70 \pm 0.045"


b.) "z= 2.57" for 99% confidence interval.


"n= 700"


"p = 0.45"


"CI = 0.45 \pm 2.57\sqrt{\dfrac{0.45 \times 0.55}{700}}"


"CI= 0.45 \pm 0.048"